# Quantum-Assisted Vehicle Routing: Realizing QAOA-based Approach on Gate-Based Quantum Computer

### v1

@@ III-A Quantum Computing
III-B Quantum Approximate Optimization Algorithm
Quantum Approximate Optimization Algorithm (QAOA) is a hybrid quantum-classical algorithm designed for combinatorial optimization problems by leveraging quantum computing properties on a near-term quantum device in the NISQ era [11]. QAOA is particularly suited for finding approximate solutions to NP-hard problems such as Max-Cut, traveling salesperson, and graph partitioning. QAOA is inspired by Quantum Annealing, which relies on the Adiabatic Theorem, which states that a system in its ground state (lowest energy) will remain there if the Hamiltonian changes slowly enough. QAOA approximates this gradual evolution using a parameterized circuit derived by Farhi et al. [9]. It works by encoding the problem into a cost Hamiltonian, then evolving a quantum state using parameterized quantum gates corresponding to the cost Hamiltonian $H_{C}$ and a mixing Hamiltonian $H_{M}$ which prepares the state
$|\psi(\gamma,\beta)\rangle=\prod_{j=1}^{p}e^{-i\beta_{j}H_{M}}e^{-i\gamma_{j}H_{C}}|+\rangle$ (2)
where |+ket|$\rangle$| + ⟩ is the even superposition state, $p$ is the number of repetitions, and $\psi$ is a parameterized state. The parameters $(\gamma,\beta)$ are optimized by a classical method based on the expected value of the cost. Once the parameters converge, the circuit can be sampled numerous times to find the solution as the most likely state. Since QAOA provides approximate solutions it is more suitable at larger scales when it becomes infeasible to solve on classical computers. While exponential speedup cannot be guaranteed, it has been shown to outperform classical heuristics in certain cases [40].
The optimization problem has to be expressed in the form of a cost Hamiltonian with no constraints. Many combinatorial problems can be naturally expressed as a quadratic unconstrained binary optimization (QUBO), which encodes optimization problems using binary variables. A key advantage of QUBO is that it can be directly mapped to the Ising Hamiltonian, a fundamental model in quantum mechanics used to describe interactions in spin systems. It allows for efficient Trotterization [41], enabling simulation of time evolution using simple quantum gates. Additionally, it makes expectation values easy to estimate via repeated measurements in the computational basis. The Ising model is particularly relevant in quantum optimization, as it forms the basis for quantum annealing making it compatible with QAOA as a cost Hamiltonian. A QUBO problem is generally expressed as:
@@ III-B Quantum Approximate Optimization Algorithm
$C(x)=\sum_{i}Q_{ii}x_{i}+\sum_{i<j}Q_{ij}x_{i}x_{j}$ (3)
To map the problem to the Ising Hamiltonian, the problem must be defined in terms of spin variables which take the values as ±$\pm 1$. This corresponds to the direction of the Z axis on the Bloch sphere, which is the computational basis. This leads to the conversion by substituting
@@
$C(z)=\sum_{i}\frac{Q_{ii}}{2}\left(z_{i}+1\right)+\sum_{i<j}\frac{Q_{ij}}{4}\left(z_{i}z_{j}+z_{i}+z_{j}+1\right)$ (4)
@@
$H_{C}=\sum_{i}h_{i}Z_{i}+\sum_{i<j}J_{ij}Z_{i}Z_{j}+\text{constant terms}$ (5)
where $Z_{i}$ is the single qubit Pauli-Z gate acting on the $i$th qubit, while $Z_{i}Z_{j}$ refers to a two qubit Pauli-Z gate defined as the tensor product
@@
$H_{M}=\sum_{i}X_{i}\,,\ \ U_{M}(\beta_{0})=e^{-i\beta_{0}H_{M}}$ (7)
The cost and mixing Hamiltonians are repeatedly applied for some $p\ll 2^{n}$ times for different values of $\gamma_{i}$ and $\beta_{i}$. The optimum values of these parameters has to be tuned using classical optimization methods, such as gradient descent.
After the optimization, the states sampled from $|\psi(\gamma,\beta)\rangle$ are likely to be the optimal solution of the optimization problem. The approximation ratio $r$ is the expected value of the cost function divided by the optimum. A single layer QAOA for Max-Cut can achieve $r\approx 0.6924$, and theoretically $r$ approaches 1 as depth increases. However, due to noise and other practical limitations deep QAOA circuits are difficult to implement, but still prove to be better than the best classical SAT solvers [26]. It has been conjectured that QAOA may be able to provide exponential speedup in certain cases [42].
III-C Vehicle Routing Problem
An instance of the capacitated VRP is characterized by a set of customers with specific demand, a fleet of vehicles with limited capacity, a central depot where vehicles start and end their routes, and a matrix representing travel costs between locations. In this study, we assume that customer demand is satisfied upon the vehicle’s arrival at the designated node. Therefore, the problem becomes to minimize the sum of the distance of all links traversed. If we denote the binary variable indicating a link between nodes $i$ and $j$, in the set of nodes $\mathcal{N}$, being active or not as $x_{i,j}$, with the distance or cost in the link as $w_{i,j}$, the objective for the link based formulation is expressed as:
$\min_{\mathbf{x}}\,\sum_{i}\sum_{j}w_{i,j}x_{i,j}\ \quad i,j\in\mathcal{N},$ (8)
@@ III-C Vehicle Routing Problem
[…]
where $S$ is any subset of the node set $\mathcal{N}$,consisting of $n$ total nodes. Meaning that any selected subset of the nodes must have at least one active link to some node in the complement set. Without these explicit constraints, subtours will appear in the solution space which seemingly satisfy all the other constraints and may have lower cost than the real solution.
@@
IV Experiment Design and Implementation
The QAOA was formulated from a QUBO representing a VRP instance based initially on a qiskit community example [43], and implemented on the 127 qubit IBM-Rensselaer quantum computer. An excerpt of the transpiled parametrized circuit is depicted in Figure 4. This parameters were optimized classically, and the final circuit was run to sample the solution.
IV-A QAOA Formulation
@@
$\begin{split}\sum_{i\sim j}w_{i,j}x_{i,j}=\ &61.323x_{0,1}+4.732x_{0,2}+61.323x_{1,0}\\ +&42.895x_{1,2}+4.732x_{2,0}+42.895x_{2,1}\end{split}$ (9)
@@
Figure 4: QAOA circuit with the left part showing the mapping to physical qubits on the quantum computer; the initialization $R_{Z}(\pi/2)$ refers to the Hadamard gate for superposition. On the right is the measurement of each qubit. The middle part has been truncated for space.
Constraints $C_{0}$ and $C_{1}$ enforce that a vehicle may leave one node only once, constraints $C_{2}$ and $C_{3}$ say that a vehicle may enter a node only once, $C_{4}$ states that there are two vehicles entering the depot, and $C_{5}$ requires that two vehicles leave the depot. Finally, $C_{6}$ eliminates sub-tours by saying that the link between nodes 1 and 2 may be utilized once or not at all, but not more than that. The statement $C_{7}$ represents that each variable $x_{i,j}$ is a binary number. The notation $i\sim j$ refers to valid connections in a graph. In this case we assume all links are valid.
The formulation was converted to a quadratic unconstrained binary optimization (QUBO) problem by creating an augmented objective function similar to the Lagrange multiplier approach [44]. This method adds to the original cost function quadratic penalties corresponding to each constraint, resulting in an unconstrained problem which has the same global optimum as the original problem. For example, $C_{0}$ can be converted to a penalty term $P(1-x_{1,0}-x_{1,2}-2x_{1,0}x_{1,2})$, where $P$ is a sufficiently large positive scalar, and $C_{6}$ may be added as $Px_{1,2}x_{2,1}$. Note that these transformations are unique to binary variables.
The specific value of $P$ should be high enough that a constraint violation results in a significant penalty as compared to the original cost function, but not too high that it hinders the optimization process. For fine-tuning purposes each constraint may even have a different penalty value to reflect its importance. As a rule of thumb, $P>\sum_{i\sim j}|w_{i,j}|$ is considered sufficient [36], but higher values are deemed more practical to aid the optimization algorithm. In this study a value of $2\sum_{i\sim j}|w_{i,j}|$ was used for all penalties.
@@
$\begin{split}&\min\mathbf{x}^{T}\mathbf{Qx}+b=\\ &875.607\,x_{0,1}x_{0,2}+875.607\,x_{0,2}x_{1,2}+875.607\,x_{1,0}x_{1,2}\\ &+875.607\,x_{1,0}x_{2,0}+875.607\,x_{0,1}x_{2,1}+218.901\,x_{1,2}x_{2,1}\\ &+875.607\,x_{2,0}x_{2,1}-1689.892\,x_{0,1}-1746.482\,x_{0,2}\\ &-1689.892\,x_{1,0}-832.712\,x_{1,2}-1746.482\,x_{2,0}\\ &-832.712\,x_{2,1}+5253.645\end{split}$ (10)
@@
$\begin{split}H_{C}&=407.14\,Z_{5}+435.43\,Z_{4}+407.14\,Z_{3}-76.17\,Z_{2}\\ &+435.43\,Z_{1}-76.17\,Z_{0}+218.9\,Z_{4}Z_{5}+218.9\,Z_{0}Z_{5}\\ &+218.9\,Z_{2}Z_{4}+218.9\,Z_{2}Z_{3}+218.9\,Z_{1}Z_{3}\\ &+54.72\,Z_{0}Z_{2}+218.9\,Z_{0}Z_{1}\end{split}$ (11)
where $[Z_{0},Z_{1},Z_{2},Z_{3},Z_{4},Z_{5}]$ respectively correspond to $[x_{0,1},x_{0,2},x_{1,0},x_{1,2},x_{2,0},x_{2,1}]$.
@@
IV-B Quantum Circuit Synthesis
@@
A cost function estimator was constructed to run the circuit with the given $\gamma$s and $\beta$s, to output the cost. Around this estimator was wrapped a classical optimizer known as Constrained Optimization by Quadratic Approximation (COBYQA) to iteratively find new values of $\gamma$ and $\beta$ which ideally bring the expected value of the cost, $\langle\psi|H|\psi\rangle$, down.
@@
Error suppression was applied to the circuit in the form of dynamic decoupling, which maintains coherence on idle qubits that can pick up unwanted interference from other qubits. This is a hardware feature that works by sending precisely timed pulses that amount to an identity operation on the qubit [45]. Pauli twirling is a technique used to turn arbitrary noise channels into specific signatures, and was also used in the experiments. The method essentially introduces Pauli rotations that eventually cancel out, increasing the fidelity of the overall circuit [46].
Figure 5: Convergence of QAOA parameters over optimization iterations
@@
The transpiled circuit was then run on the IBM-Rennselaer Eagle r3 quantum computer, which took around 14 minutes. After these values converged as depicted in Figure 5, the final parameters were assigned and the circuit executed again for 10,000 shots, which take about 15 seconds.
IV-C Results and Interpretation
@@
The results obtained from running the QAOA quantum circuit for the VRP provide insights into the effectiveness and scalability of quantum computing for combinatorial optimization tasks. The final sampling using the trained parameters resulted in a distribution of results, shown in Figure 6. It can be seen that the most likely bit-string is 111010, which corresponds to $x_{0,1},x_{0,2},x_{1,0},x_{2,0}$ having the value 1, and the remaining variables are 0.
Figure 6: Sampled distribution from optimized circuit
@@
This outcome suggests that the algorithm was able to identify the optimal solution with a high probability. However, the probability distribution also shows that suboptimal solutions were sampled with non-negligible frequencies, a behavior that could be attributed to noise in the quantum hardware, low depth $p$, and limitations in the standard QAOA parameter tuning process.
Figure 7: Assigned routes for the 3 node, 2 vehicle problem
@@
The routing solution corresponding to the most probable bit-string is shown in Figure 7. The selected routes by QAOA successfully satisfy the constraints of the 3-node, 2 vehicle problem, demonstrating that QAOA was able to generate a feasible routing configuration.
@@
Our findings also highlight the rapid growth in quantum gate counts as the problem scales. Larger instances that require more depth quickly become too large to run practically. For example a 4 node problem with depth of 2 generates a 12 qubit circuit consisting of 1685 $R_{Z}$, 883 $R_{X}$, 425 echoed cross-resonance, and 85 Pauli-X gates. This problem took over 4 hours to run on the quantum computer and generated an infeasible solution. In fact, the top 10 most sampled states were all infeasible. Similarly, a 5 node problem with depth 2 transpiled into a 20 qubit circuit of about 5000 basic gates. The number of circuit gates increases linearly with $p$ as plotted in Figure 8(a), while the total gates increases exponentially against the number of nodes in the problem. However, a more representative measure of circuit depth from a noise accumulation perspective is the maximum number of two qubit gates in a path from the start of the circuit to the end. This metric also exhibits an exponential increase as shown in Figure 8(b).
@@
(b) Circuit depth scaling with $n$, for different values of $p$. The spread around the lines indicates variance for different instances of the problem.
Figure 8: Total circuit gates and circuit depth for different versions of the problem
IV-D Circuit complexity

### v2

@@ 3.1. Quantum Computing
3.2. Quantum Approximate Optimization Algorithm
Quantum Approximate Optimization Algorithm (QAOA) is a hybrid quantum-classical algorithm designed for combinatorial optimization problems by leveraging quantum computing properties on a near-term quantum device in the NISQ era (Farhi and Harrow, 2016). QAOA is particularly suited for finding approximate solutions to NP-hard problems such as Max-Cut, traveling salesperson, and graph partitioning. QAOA is inspired by Quantum Annealing, which relies on the Adiabatic Theorem, which states that a system in its ground state (lowest energy) will remain there if the Hamiltonian changes slowly enough. QAOA approximates this gradual evolution using a parameterized circuit derived by Farhi et al. (Farhi et al., 2014). It works by encoding the problem into a cost Hamiltonian, then evolving a quantum state using parameterized quantum gates corresponding to the cost Hamiltonian $H_{C}$ and a mixing Hamiltonian $H_{M}$ which prepares the state
(2) $|\psi(\gamma,\beta)\rangle=\prod_{j=1}^{p}e^{-i\beta_{j}H_{M}}e^{-i\gamma_{j}H_{C}}|+\rangle$
where $|+\rangle$ is the equal superposition state (a common choice for the initial state which must be the ground state of the mixing Hamiltonian), $p$ is the number of repetitions, and $\psi$ is a parameterized state. The parameters $(\gamma,\beta)$ are optimized by a classical method based on the expected value of the cost. Once the parameters converge, the circuit can be sampled numerous times to find the solution as the most likely state. Since QAOA provides approximate solutions it is more suitable at larger scales when it becomes infeasible to solve on classical computers. While exponential speedup cannot be guaranteed, it has been shown to outperform classical heuristics in certain cases (Zhou et al., 2020).
The optimization problem has to be expressed in the form of a cost Hamiltonian with no constraints. Many combinatorial problems can be naturally expressed as a quadratic unconstrained binary optimization (QUBO), which encodes optimization problems using binary variables. A key advantage of QUBO is that it can be directly mapped to the Ising Hamiltonian, a fundamental model in quantum mechanics used to describe interactions in spin systems. It allows for efficient Trotterization (Kluber, 2023), enabling simulation of time evolution using simple quantum gates. Additionally, it makes expectation values easy to estimate via repeated measurements in the computational basis. The Ising model is particularly relevant in quantum optimization, as it forms the basis for quantum annealing making it compatible with QAOA as a cost Hamiltonian. A QUBO problem is generally expressed as:
@@ 3.2. Quantum Approximate Optimization Algorithm
(3) $C(\mathbf{x})=\sum_{i}Q_{ii}x_{i}+\sum_{i<j}Q_{ij}x_{i}x_{j}$
To map the problem to the Ising Hamiltonian, the problem must be defined in terms of spin variables which take the values as $\pm 1$. This corresponds to the direction of the Z axis on the Bloch sphere, which is the computational basis. This leads to the conversion by substituting
@@
(4) $C(\mathbf{z})=\sum_{i}\frac{Q_{ii}}{2}\left(z_{i}+1\right)+\sum_{i<j}\frac{Q_{ij}}{4}\left(z_{i}z_{j}+z_{i}+z_{j}+1\right)$
@@
(5) $C(\mathbf{x})=C(\mathbf{z})=\langle\mathbf{x}|H_{C}|\mathbf{x}\rangle$
where $H_{C}$ is the cost Hamiltonian operator
(6) $H_{C}=\sum_{i}h_{i}Z_{i}+\sum_{i<j}J_{ij}Z_{i}Z_{j}+\text{constant terms}$
for some real coefficients $h_{i}$, and $J_{ij}$, where $Z_{i}$ is the single qubit Pauli-Z gate acting on the $i$th qubit, while $Z_{i}Z_{j}$ refers to a two qubit Pauli-Z gate defined as the tensor product
@@
(8) $H_{M}=\sum_{i}X_{i}\,,\ \ U_{M}(\beta_{0})=e^{-i\beta_{0}H_{M}}$
The choice of $Z$ and $ZZ$ for the cost and $X$ for the mixer is due to the requirement that both Hamiltonians do not commute. The cost and mixing Hamiltonians are repeatedly applied for some $p\ll 2^{n}$ times for different values of $\gamma_{i}$ and $\beta_{i}$. The optimum values of these parameters has to be tuned using classical optimization methods, such as gradient descent.
After the optimization, the states sampled from $|\psi(\gamma,\beta)\rangle$ are likely to approximate the optimal solution of the optimization problem. The approximation ratio $r$ is the expected value of the cost function divided by the optimum. A single layer QAOA for Max-Cut can achieve $r\approx 0.6924$, and theoretically $r$ approaches 1 as depth increases. However, due to noise and other practical limitations deep QAOA circuits are difficult to implement, but still prove to be better than the best classical SAT solvers (Boulebnane and Montanaro, 2024). It has been conjectured that QAOA may be able to provide exponential speedup in certain cases (Montanez-Barrera and Michielsen, 2024).
3.3. Vehicle Routing Problem
An instance of the capacitated VRP is characterized by a set of customers with specific demand, a fleet of vehicles with limited capacity, a central depot where vehicles start and end their routes, and a matrix representing travel costs between locations. In this study, we assume that customer demand is satisfied upon the vehicle’s arrival at the designated node. Therefore, the problem becomes to minimize the sum of the distance of all links traversed. If we denote the binary variable indicating a directional link between nodes $i$ and $j$, in the set of nodes $\mathcal{N}$, being active or not as $x_{i,j}$, with the distance or cost in the link as $w_{i,j}$, the objective for the link based formulation is expressed as:
(9) $\min_{\mathbf{x}}\,\sum_{i}\sum_{j}w_{i,j}x_{i,j}\ \quad i,j\in\mathcal{N},$
@@ 3.3. Vehicle Routing Problem
$\sum_{i\in S}\sum_{j\notin S}x_{i,j}\geq 1,\quad\forall S\subset\mathcal{N},\ 2\leq|S|\leq n-1$
where $S$ is any subset of the node set $\mathcal{N}$, consisting of $n$ total nodes. Meaning that any selected subset of the nodes must have at least one active edge to some node in the complement set. Without these explicit constraints, subtours will appear in the solution space which seemingly satisfy all the other constraints and may have lower cost than the real solution. For example, in an 8-node, 2-vehicle problem there could be the set of routes $0\rightarrow 1\rightarrow 2\rightarrow 0$, $0\rightarrow 3\rightarrow 4\rightarrow 0$, and a subtour $5\rightarrow 6\rightarrow 7\rightarrow 5$. This set satisfies all constraints except subtour elimination.
@@
Another encoding method uses the time step expansion by including time steps into the decision variable $x^{k}_{v.t}$, which is 1 if the $k$th vehicle visits the node $v$ at time step $t$. The optimization problem is then
$\min_{\mathbf{x}}\sum_{v,v^{\prime},t,k}w_{vv^{\prime}}\cdot x_{v,t}^{k}\cdot x_{v^{\prime},t+1}^{k}$
s.t. $\sum_{v=1}^{n}\left(1-\sum_{k=1}^{K}\sum_{t=1}^{T}x_{v,t}^{k}\right)^{2}=0$
$\sum_{k=1}^{K}\sum_{t=1}^{T}\left(1-\sum_{v=0}^{n}x_{v,t}^{k}\right)^{2}=0$
$\sum_{t=1}^{T}\sum_{v=1}^{n}x_{v,t}^{k}\cdot d_{v}\leq C\quad\forall k\in K$
where $d_{v}$ is the demand of a node and $C$ is the capacity of a vehicle. The assumption of demands always being met simplifies the last constraint, but the qubit requirements for this formulation are $O(n\cdot T\cdot K)$. This method was used in (Palackal et al., 2023) with a clustering preprocessing phase to simplify the problem.
4. Experiment Design and Implementation
The QAOA was formulated from a QUBO representing a VRP instance based initially on a qiskit community example (Community, 2024), and implemented on the 127 qubit IBM-Rensselaer quantum computer. An excerpt of the transpiled parametrized circuit is depicted in Figure 4. This parameters were optimized classically, and the final circuit was run to sample the solution.
4.1. QAOA Formulation
@@
(10) $\begin{split}\sum_{i\sim j}w_{i,j}x_{i,j}=\ &61.323x_{0,1}+4.732x_{0,2}+61.323x_{1,0}\\ +&42.895x_{1,2}+4.732x_{2,0}+42.895x_{2,1}\end{split}$
@@
Figure 4. QAOA circuit with the left part showing the mapping to physical qubits on the quantum computer; the initialization $R_{Z}(\pi/2)$ refers to the Hadamard gate for superposition. On the right is the measurement of each qubit. The middle part has been truncated for space.
Constraints $C_{0}$ and $C_{1}$ enforce that a vehicle may leave one node only once, constraints $C_{2}$ and $C_{3}$ say that a vehicle may enter a node only once, $C_{4}$ states that there are two vehicles entering the depot, and $C_{5}$ requires that two vehicles leave the depot. Finally, $C_{6}$ eliminates sub-tours by saying that the link between nodes 1 and 2 may be utilized once or not at all, but not more than that. This removes the possibility of a vehicle going from node 1 to 2 then back 2 to 1. The statement $C_{7}$ represents that each variable $x_{i,j}$ is a binary number. The notation $i\sim j$ refers to valid connections in a graph. In this case we assume all links are valid.
The formulation was converted to a quadratic unconstrained binary optimization (QUBO) problem by creating an augmented objective function similar to the Lagrange multiplier approach (Glover et al., 2018). This method adds to the original cost function quadratic penalties corresponding to each constraint, resulting in an unconstrained problem which has the same global optimum as the original problem. For example, $C_{0}$ can be converted to a penalty term $P(1-x_{1,0}-x_{1,2}+2x_{1,0}x_{1,2})$, where $P$ is a sufficiently large positive scalar, and $C_{6}$ may be added as $Px_{1,2}x_{2,1}$. Note that these transformations are unique to binary variables.
The specific value of $P$ should be high enough that a constraint violation results in a significant penalty as compared to the original cost function, but not too high that it hinders the optimization process. For fine-tuning purposes each constraint may even have a different penalty value to reflect its importance. As a rule of thumb, $P>\sum_{i\sim j}|w_{i,j}|$ is considered sufficient (Harwood et al., 2021), but higher values are deemed more practical to aid the optimization algorithm. In this study a value of $2\sum_{i\sim j}|w_{i,j}|$ was used for all penalties.
@@
(11) $\begin{split}&\min\mathbf{x}^{T}\mathbf{Qx}+b=\\ &875.607\,x_{0,1}x_{0,2}+875.607\,x_{0,2}x_{1,2}+875.607\,x_{1,0}x_{1,2}\\ &+875.607\,x_{1,0}x_{2,0}+875.607\,x_{0,1}x_{2,1}+218.901\,x_{1,2}x_{2,1}\\ &+875.607\,x_{2,0}x_{2,1}-1689.892\,x_{0,1}-1746.482\,x_{0,2}\\ &-1689.892\,x_{1,0}-832.712\,x_{1,2}-1746.482\,x_{2,0}\\ &-832.712\,x_{2,1}+5253.645\end{split}$
@@
(12) $\begin{split}H_{C}&=407.14\,Z_{5}+435.43\,Z_{4}+407.14\,Z_{3}-76.17\,Z_{2}\\ &+435.43\,Z_{1}-76.17\,Z_{0}+218.9\,Z_{4}Z_{5}+218.9\,Z_{0}Z_{5}\\ &+218.9\,Z_{2}Z_{4}+218.9\,Z_{2}Z_{3}+218.9\,Z_{1}Z_{3}\\ &+54.72\,Z_{0}Z_{2}+218.9\,Z_{0}Z_{1}\end{split}$
where $Z_{0},Z_{1},Z_{2},Z_{3},Z_{4},Z_{5}]$ respectively correspond to $x_{0,1},x_{0,2},x_{1,0},x_{1,2},x_{2,0},x_{2,1}]$.
@@
The mixer Hamiltonian, $H_{M}$, is chosen to be the commonly used $H_{M}=X^{\otimes n}$, where $X$ is the Pauli-$X$ gate.
4.2. Quantum Circuit Synthesis
@@
[…]
@@
Error suppression was applied to the circuit in the form of dynamic decoupling, which maintains coherence on idle qubits that can pick up unwanted interference from other qubits. This is a hardware feature that works by sending precisely timed pulses that amount to an identity operation on the qubit (Ezzell et al., 2023). Pauli twirling is a technique used to turn arbitrary noise channels into specific signatures, and was also used in the experiments. The method essentially introduces Pauli rotations that eventually cancel out, increasing the fidelity of the overall circuit (Wallman and Emerson, 2016).
Figure 5. Convergence of QAOA parameters over optimization iterations
@@
The transpiled circuit was then used to run the hybrid QAOA on the IBM-Rennselaer Eagle r3 quantum computer, and the optimization process took around 14 minutes. After these values converged as depicted in Figure 5, the final parameters were assigned and the circuit executed again for 10,000 shots, which take about 15 seconds of wall clock time.
4.3. Results and Interpretation
@@
The results obtained from running the hybrid QAOA algorithm for the VRP provide insights into the effectiveness and scalability of quantum computing for combinatorial optimization tasks. The final sampling using the trained parameters resulted in a distribution of results, shown in Figure 6. It can be seen that the most likely bit-string is 111010, which corresponds to $x_{0,1},x_{0,2},x_{1,0},x_{2,0}$ having the value 1, and the remaining variables are 0.
Figure 6. Sampled distribution from optimized circuit
Figure 7. Assigned routes for the 3 node, 2 vehicle problem. Each arrow head indicates a directed edge used in the solution, and the red numbers are node labels.
@@
This outcome suggests that the algorithm was able to identify the optimal solution with a high probability. However, the probability distribution also shows that suboptimal solutions were sampled with non-negligible frequencies, a behavior that could be attributed to noise in the quantum hardware, low depth $p$, and limitations of the standard QAOA. For finite $p$ QAOA is not expected to return the optimal solution with probability 1.
@@
The routing solution corresponding to the most probable bit-string is shown in Figure 7. The three nodes are labeled in red, with 0 being the depot node. The selected routes (depicted by arrows) from QAOA successfully satisfy the constraints of the 3-node, 2 vehicle problem, demonstrating that QAOA was able to generate a feasible routing configuration.
@@
Our findings also highlight the rapid growth in quantum gate counts as the problem scales. Larger instances that require more depth quickly become too large to run practically. For example a 4 node problem with depth of 2 generates a 14 qubit circuit consisting of 3207 $R_{Z}$, 1642 $R_{X}$, 928 echoed cross-resonance, and 214 Pauli-X gates. This problem took over 2 hours to run on the quantum computer and generated an infeasible solution. In fact, the top 10 most sampled states were all infeasible. Similarly, a 5 node problem with depth 2 transpiled into a 30 qubit circuit of about 30,000 basic gates. The number of circuit gates increases linearly with $p$ as plotted in Figure 8(a), while the total gates increases exponentially against the number of nodes in the problem. However, a more representative measure of circuit depth from a noise accumulation perspective is the maximum number of two qubit gates in a path from the start of the circuit to the end. This metric exhibits an exponential increase as shown in Figure 8(b).
@@
(b) Circuit depth scaling with $n$, for different values of pp. The shading indicates variance for different instances of the problem.
Figure 8. Total circuit gates and circuit depth for different versions of the problem
4.4. Circuit complexity